# A Universal Relation Between Intermittency and Dissipation in Turbulence

### v2

@@ Abstract
The relationship between fundamental quantities of turbulent flows like the dissipation constant $C_{\varepsilon}$ and the intermittency factor $\mu$, expected to be constant for homogeneous isotropic turbulence, is explored with respect to their relationship for a wider class of non-ideal turbulent flows. For the energy cascade, $C_{\varepsilon}$ is known to reflect its basic overall properties, while $\mu$ quantifies the emerging intermittency down the cascade. Using a very large dataset of turbulent wakes, grid-generated turbulence and an axisymmetric jet, we find that $\mu$ is inversely proportional to $C_{\varepsilon}$, providing a new empirical principle.
@@
Turbulence is still one of the major unsolved problems in physics [1]. The lack of a theory of turbulent flows hinders our modelling capabilities in several fields, as there are many situations where turbulence has a crucial influence, with turbulent wakes generated by wind turbines being currently among the most relevant industry-related applications [2, 3]. The most common approach to tackle this problem is to define equations where factors or exponents are obtained by dimensional arguments or experiments. Within this approach, a relevant question that arises concerns the universality of these constants [4, 5]. Since the publication of Kolmogorov’s phenomenology in 1941 (K41) [6], there has been intense work to shed light on this question. This was enhanced by the publication of Kolmogorov’s revised theory in 1962 [7], as further constants appeared. Although there were some deviations in experimental results, the overall opinion – at least until the 1990s – was that such constants, for sufficiently large Reynolds numbers, are universal for fully developed turbulence, which led to a focus on their numerical values [4, 8, 9, 10, 11].
@@
Among all relevant parameters, two of the most important constants to describe turbulent flows are, arguably, the dissipation constant $C_{\varepsilon}$ and the intermittency factor $\mu$. The former has been identified as one of the most important parameters to describe turbulence [12], as it can be directly linked to the energy cascade [13]. On the other hand, the constant $\mu$, first introduced by Kolmogorov’s 1962 theory [7], and later redefined by Castaing and collaborators [14, 9], quantifies departures from self-similarity within the inertial range. Despite its relevance, a physical interpretation of $\mu$ is still missing. Furthermore, over the last two decades, the universality of $C_{\varepsilon}$ has been challenged by several experimental and numerical studies [13].
@@
with $\varepsilon$ the mean turbulent kinetic energy dissipation rate, $u^{\prime}$ the RMS value of the streamwise fluctuating velocity and $L$ the integral length scale. The importance of $C_{\varepsilon}$ for modelling arises from the fact that it gives an estimation of the mean dissipation rate, which is often needed to have a closed system of equations [15]. According to K41 phenomenology, considering stationary, homogeneous turbulence, $C_{\varepsilon}$ should depend on boundary conditions, being constant for sufficiently large values of the Taylor-length based Reynolds number $Re_{\lambda}=u^{\prime}\,\lambda/\nu$, with $\lambda$ the Taylor length scale and $\nu$ the kinematic viscosity of the fluid.
@@
In several free-shear flows, results not consistent with K41 phenomenology have been reported, and the value of $C_{\varepsilon}$ at the centreline has been found to decrease with increasing $Re_{\lambda}$ [13, 16]. More precisely, for this so-called non-equilibrium cascade, $C_{\varepsilon}$ is expected to be linear with $\sqrt{Re_{G}}/Re_{\lambda}$ (with $Re_{G}$ a Reynolds number that depends on the inlet properties of the flow, such as the wind tunnel inflow velocity and the diameter of the wake-generating object). Most of the work on this topic aims at quantifying the dependence of $C_{\varepsilon}$ over a broad range of Reynolds numbers. These studies concern laminar inflows, and little attention has been paid to further details of the energy cascade, such as emerging intermittency.
The intermittency factor $\mu$ is used to characterise deviations from K41 and is a measure of how “fast” the cascade propagates to smaller scales. This deviation is seen in the changing statistics of the averaged velocity increments $\delta_{u}(r):=u(x+r)-u(x)$ ($r$ being the spatial increments), which can be quantified using the scale-dependent shape factor $\Lambda^{2}(r)$. $\Lambda^{2}$, that is linked to the dispersion of the logarithm of the – also scale-dependent – energy dissipation [7], can be estimated as [17],
$\Lambda^{2}(r)=\frac{\ln\left(F(\delta_{u}(r))/3\right)}{4},$ (2)
@@
where $F(\delta_{u}(r))$ denotes the flatness. In the inertial range, for $\eta\ll r\ll L$ (with $\eta=(\nu^{3}/\varepsilon)^{\frac{1}{4}}$, being the Kolmogorov length scale), the scale dependence is expected to follow the relation [7, 14],
${\Lambda^{2}=\Lambda_{0}^{2}+\mu/9\;\ln(L/r),}$ (3)
@@
with $\Lambda_{0}^{2}$ a constant that depends on the properties of the flow (see also figure 1). Note, that the relation (3) is consistent with the power-law scaling behavior of the energy spectral density in the inertial range, as discussed below (see equation 4).
@@
Within this rationale, the non-Gaussianity of the velocity increment distributions can be quantified by a single scalar parameter. The value of $\mu$ has been commonly accepted to be constant at large values of the $Re_{\lambda}$. The reported values range between $0.2$ and $0.4$ for homogeneous isotropic turbulence (HIT) and a value of $0.26$ has been established as representative of turbulent flows [9, 18, 3].
Despite intense research on turbulent flows, no relation between $\mu$ and $C_{\varepsilon}$ has been predicted nor found. Furthermore, the universality of these constants remains an open question, with a growing amount of evidence pointing against this feature. While most studies deal with homogeneous turbulence, in this study we also focus on inhomogeneous flows. Hence, we present a detailed experimental work on turbulent wakes supported by previously published datasets on grid-generated turbulence and a turbulent axisymmetric jet.
@@
Figure 1: Schematic representation of the spatial expansion of the turbulence cascade: the blue curve shows the energy spectrum, with energy transported to smaller scales, while the red curve illustrates the evolution of the shape factor $\Lambda^{2}(r)$ for normalised velocity increments.
@@
We carried out wind tunnel experiments in LEGI, Grenoble by generating a turbulent wake behind porous and impermeable objects under different inflow conditions. The facility is a closed-circuit system with test section dimensions of 0.75 $\times$ 0.75 $\times$ 444\,4 m3 (see supplementary material [19]). The objects were cylinders and disks while both the diameter and solidity vary (see also table I in the supplementary material [19]). To reduce the blockage, the cylinders were directly screwed to the walls of the tunnel while the disks were hold at the tunnel’s centreline with four thin piano wires which went through taped holes in the tunnel walls. For varying the inflow, both the entrance of the test section and the tunnel freestream velocity were modified. We used laminar inflow, and different turbulent background flows generated by a static regular grid with cylindrical bars or an active grid driven in a triple random mode (information about the turbulent flow produced by grids is reported in a previous work [20]). The wake of the objects was scanned by a 111\,1D-hot-wire with a temporal resolution of 505050\,50kHz (with an anti-aliasing filter set at 303030\,30kHz) which was traversed in streamwise and spanwise directions. The conversion from the temporal to the spatial domain was made using Taylor’s hypothesis. The nearest streamwise position to the wake generator was $5\,D$ and the farthest was $200\,D$, with $D$ being the diameter of the object. For every position the data was acquired for 120120120\,120s. In total, $15$ configurations were carried out and $505$ centreline points were taken. For simplicity, only points in centreline are considered 111pre-analysis shows that our results also hold for non-centreline data. Additionally, the background flows without the object were also measured (see table I in the supplementary material [19] for further information.)
@@
As discussed above, we complement our study with $113$ velocity time series from three previously published experimental datasets covering different canonical flows [22, 23, 24] (for further details see supplementary material [19]). Within those measurement configurations, five involve decaying grid-generated turbulence and one captures a free axisymmetric jet. All cases use air as the fluid and the velocity time signals were collected by means of a 111\,1D-hot-wire.
@@
For all those centreline points from all mentioned datasets, a rigorous analysis of common 111\,1D-free-shear turbulence constants is derived and checked individually. Only time series that have the following characteristics were kept for further analysis:
@@
a) the turbulence intensity $TI=u^{\prime}/\overline{u}$ with $\overline{u}$ being the actual mean velocity, decreases in streamwise direction, b) the large scale increments have an almost Gaussian distribution ($\Lambda_{0}^{2}<0.025$) c) the slope of the energy spectrum within the inertial range is smaller than $-1.25$, d) the length of the inertial range for the energy spectrum is larger than $0.5$ decades, e) the $R^{2}$-value of the linear regression of the fit of $\Lambda^{2}$ is larger than $0.99$, f) the mean squared deviation from the calculation of $\Lambda^{2}$ is smaller than $0.11$, g) the flatness of the velocity time signal is around $3$ ($3.12.67<F<$ 3.1) and h) the skewness of the time signal is below $0.1$. Details and an example of the analysis is shown and discussed in the supplementary material [19]. After this conditioning, 313 velocity time series remained for the analysis. Each of them represents at least $2\times 10^{4}$ integral time scales.
@@
The integral length scale $L$ is calculated by integrating the autocorrelation function until it crosses $1/\mathrm{e}$ ($\mathrm{e}$ being the Eulerian number). Other integration limits were tested, and the trends presented in this Letter remained unchanged. $\varepsilon$ is derived – the common definition for HIT conditions is used – by integrating the dissipation spectrum $\varepsilon=\int_{0}^{\infty}15\nu\>k^{2}E(k)\>\mathrm{d}k$ (with $k$ the wavenumber in the Fourier space) and modelling the high-frequency noise with an exponential fit (a test with different fitting limits and functions gave an estimated error in the estimation of $\varepsilon$ in the order of 5%). We extracted, after equation (3), $\mu$ by fitting the evolution of the shape factor $\Lambda^{2}$ within the inertial range [14].
@@
First, we investigate if the parameters $\mu$ and $C_{\varepsilon}$ present a constant value in our dataset, including all the flows studied. In figure 2a) it is clear that $\mu$ presents a large spreading and it does not approach towards an asymptotic limit at large values of $Re_{\lambda}$. (For reference, the dependency of $\mu$ with the streamwise distance to the turbulence generator is shown in the supplementary material [19].) Figure 2b) shows that $C_{\varepsilon}$ also behaves similarly, i.e. $C_{\varepsilon}$ does not have a constant value.
@@
Having a closer look on the data we see that for each set of boundary conditions, a linear relation between $C_{\varepsilon}$ and $\sqrt{Re_{G}}/Re_{\lambda}$ is consistent with the dissipation scalings from non-equilibrium turbulence [13]. Remarkably, when such conditions are changed, the linear relation still applies but with a different slope, showing the emergence of new length scales that affect the dissipation scalings.
@@
Overall, our dataset does not show a universal or predictable behaviour in terms of neither $C_{\varepsilon}$ nor $\mu$. Next, we want to see if there is a relation between $\mu$ and $C_{\varepsilon}$. Figure 3a) shows that $\mu\propto 1/C_{\varepsilon}$, or, respectively, that the product of those two quantities is constant. In order to quantify this, initially a simple fit is presented through the data points. Note that the commonly accepted value of $\mu$ of around 0.26 is not the asymptotic limit of our measurements. Moreover, the product of $\mu$ and $C_{\varepsilon}$, here defined as $\alpha$, is shown as a function of $Re_{\lambda}$ in figure 3b). Remarkably, $\alpha$ not only remains approximately constant for all cases, it also does not show any trends with $Re_{\lambda}$. In the supplementary material we provide further figures and a table showing that the RMS value of $\alpha$ remains lower than the individual corresponding RMS values of $C_{\varepsilon}$ and $\mu$ [19]. To the authors best knowledge, this is the first time a clear relation between these constants is found.
[…]
@@
Figure 3: a) $\mu$ as a function of $C_{\varepsilon}$ for the highly restricted data. The red line indicates a commonly accepted value for $\mu$ for homogeneous isotropic turbulence [9], and the black line corresponds to a least-square fit ($R^{2}=0.86$ with the fitted values of $0.1067\pm 0.0025$ and $0.0227\pm 0.0091$). b) $\alpha$ versus $Re_{\lambda}$ for the highly restricted data. The red line indicates the result for $\alpha$ from the fit from a) while the black solid line represents the actual mean value of the ensemble of $\alpha$ values. Additionally, two black dashed lines indicate the corresponding standard deviations from the mean. For both figures, the symbols and corresponding configurations are shown and explained in figure 2 and table I in the supplementary material [19].
@@
We continue by investigating how two other relevant parameters behave, namely the Kolmogorov constant $C_{k}$ and the power law governing the energy spectral density within the inertial range. Following the notation of  [25],
${E(k)=C_{k}\>\>\varepsilon^{2/3}\>k^{-5/3}\>(k\>\eta)^{-\gamma+5/3},}$ (4)
[…]
@@
In figure 4a) & b), the dependency of $\gamma$ on $C_{\varepsilon}$ as well as $C_{k}$ on $\gamma$ are presented, respectively ($C_{k}$ as a function of $C_{\varepsilon}$ is shown in the supplementary material [19]). Interestingly, the relation between $\mu$ and $C_{\varepsilon}$ is transferred to the other quantities. Moreover, all quantities present large variations with respect to the expected values for HIT within the K41 phenomenology. We remark that our results are still consistent with those reported in other experiments in inhomogeneous flows, such as other turbulent wakes [3] or atmospheric turbulence [27].
@@
Figure 4: a) $\gamma$ as a function of $C_{\varepsilon}$ and b) $C_{k}$ versus $\gamma$ for the highly restricted data. The red lines indicate both the commonly accepted value for $C_{k}$ and $\gamma$ for HIT [8], while the black lines correspond to a least-square fit for a) and b) with $R^{2}$ being $0.8$ and $0.98$, respectively. Once more, the symbols and corresponding configurations are shown and explained in figure 2 and table I in the supplementary material [19].

### v3

@@ Abstract
Fundamental quantities of turbulent flows, such as the dissipation constant $C_{\varepsilon}$ and the intermittency factor $\mu$, are examined in relation to each other for a broader class of non-ideal turbulent flows. In the context of the energy cascade, it is known that $C_{\varepsilon}$ reflects its basic overall properties, while $\mu$ quantifies the intermittency that emerges throughout the cascade. Using an extensive hot-wire dataset of turbulent wakes, grid-generated turbulence, and an axisymmetric jet, we individually analyze these quantities as one-dimensional surrogates of the energy cascade, considering only data that exhibit consistent scaling behavior. We find that $\mu$ is inversely proportional to $C_{\varepsilon}$, offering a new empirical principle that bridges the gap between large and small scales in arbitrary turbulent flows.
preprint: AIP/123-QED
Despite recent advances in the study of turbulent flows sreenivasan2025turbulence, solving the Navier–Stokes equations at large Reynolds numbers and for an arbitrary set of boundary conditions still remains beyond current numerical and theoretical capabilities frisch1995turbulence; alberti2023still. The lack of a simplifying framework for complex turbulent flows hinders modeling efforts across various fields, as there are many engineering problems where turbulence plays a crucial role, turbulent wakes generated by wind turbines being among the most relevant industry-related applications stevens2017flow; neunaber2020distinct. The most common approach to tackle turbulence is to define equations in which factors or exponents are determined by dimensional arguments or experiments. Within this approach, a relevant question that arises concerns the universality of these constants, even when considering homogenous isotropic turbulence (HIT) only praskovsky1994measurements; renner2002universality.
@@
Since the publication of Kolmogorov’s phenomenology in 1941 (K41) kolmogorov1941local, there has been intense work to shed light on this question. This was enhanced by the publication of Kolmogorov’s revised theory in 1962 kolmogorov1962refinement and further developments like the multifractal formalism (cf. sreenivasan1991fractals), as further constants appeared. Although there were some deviations in experimental results, the overall opinion – at least until the 1990s – was that such constants, for sufficiently large Reynolds numbers, are universal for fully developed turbulence, which led to a focus on their numerical values praskovsky1994measurements; sreenivasan1995universality; arneodo1996structure; praskovsky1997comprehensive; yeung1997universality.
@@
Among all relevant parameters, two of the most important constants to describe turbulent flows are, arguably, the dissipation constant $C_{\varepsilon}$ and the intermittency factor $\mu$. The former has been identified as one of the most important parameters to describe turbulence lumley1992some, as it can be directly linked to the energy cascade vassilicos2015dissipation. On the other hand, the constant $\mu$, first introduced by Kolmogorov’s 1962 theory kolmogorov1962refinement, and later redefined by Castaing and collaborators castaing1990velocity; arneodo1996structure, quantifies departures from self-similarity within the inertial range. Despite its relevance, a physical interpretation of $\mu$ is still missing. Nevertheless, and independently on the formalism used to interpret this quantity, it provides a scalar value that quantifies the relevance of intermittency as departures from Gaussianity of velocity derivatives and their increments in the inertial range. Furthermore, over the last two decades, the universality of $C_{\varepsilon}$ (sometimes referred to as “zeroth law” of turbulence sreenivasan2025turbulence) has been challenged by several experimental and numerical studies vassilicos2015dissipation.
@@
with $\varepsilon$ the mean turbulent kinetic energy dissipation rate, $u^{\prime}$ the RMS value of the streamwise fluctuating velocity and $L$ the integral length scale. The importance of $C_{\varepsilon}$ for modelling arises from the fact that it gives an estimation of the mean dissipation rate, which is often needed to have a closed system of equations george1989self. According to K41 phenomenology, considering stationary, homogeneous turbulence, $C_{\varepsilon}$ should depend on boundary conditions, being constant for sufficiently large values of the Taylor-length based Reynolds number $Re_{\lambda}=u^{\prime}\,\lambda/\nu$, with $\lambda$ the Taylor length scale and $\nu$ the kinematic viscosity of the fluid.
@@
In several free-shear flows, results not consistent with K41 phenomenology have been reported, and the value of $C_{\varepsilon}$ at the centreline has been found to decrease with increasing $Re_{\lambda}$ vassilicos2015dissipation; ortiz2021high. More precisely, for this so-called non-equilibrium cascade, $C_{\varepsilon}$ is expected to be linear with $\sqrt{Re_{G}}/Re_{\lambda}$ (with $Re_{G}$ a Reynolds number that depends on the inlet properties of the flow, such as the wind tunnel inflow velocity and the diameter of the wake-generating object). Most of the work on this topic aims at quantifying the dependence of $C_{\varepsilon}$ over a broad range of Reynolds numbers. These studies concern laminar inflows, and little attention has been paid to further details of the energy cascade, such as emerging intermittency.
Intermittency is one of the causes of deviations from the K41 phenomenology. One of its most significant features is that velocity increments, $u_{r}(x):=u(x+r)-u(x)$, exhibit extreme events at small scales, $r$. There are two standard approaches to quantify intermittency.
The most common one is to investigate the scaling behavior of generalized structure functions of a given order $n$, i.e., the $n$-th order moment of velocity increments, $\overline{(u_{r}(x))^{n}}$, where the over-bar implies a spatial averaging. As K41 implies that $\overline{(u_{r}(x))^{n}}\propto(\varepsilon\>r)^{\zeta_{n}}$, with the exponent $\zeta_{n}$ equal to $n/3$, many works focus on corrections either to the power-law form or to the value of the exponent sinhuber2017dissipative; reinke2018universal, cf. sreenivasan1997phenomenology. This approach has been extensively used for HIT, resulting in various scaling models for the exponent $\zeta_{n}$ of the structure functions. Starting from Kolmogorov’s 1962 theory (based on a lognormal model for the two-point energy dissipation), many improvements and alternative frameworks have been proposed, such as the works of She-Leveque she1994universal, Castaing-Yakhot castaing1996temperature; yakhot1998probability, and L’vov-Procaccia l2000analytic. For a recent review, the reader is referred to Benzi & Toschi benzi2023lectures.
In this structure function approach, two questions arise. On the one hand, as discussed above, an open question is how the higher-order exponents $\zeta_{n}$ deviate from the relation $\zeta_{n}=n/3$ implied by K41. Typically, these deviations become significant for $n>4$, while differences between anomalous scaling models become significant for $n>6$. On the other hand, another open discussion concerns how deviations from the expected scaling of the structure function can be understood in the context of non-isotropic turbulent flows Arad1998; biferale2002probing; kurien2002measures; benzi2023lectures; zhou2021turbulence.
A second approach is to investigate directly the form of the probability density functions $p(u_{r}(x))$ at different values of $r$. This can be done via the scale-dependent shape factor $\Lambda^{2}(r)$. This quantity, that is linked to the dispersion of the logarithm of the – also scale-dependent – energy dissipation kolmogorov1962refinement; castaing1990velocity, can be estimated as chilla1996multiplicative,
$\Lambda^{2}(r)=\frac{\ln\left(F(u_{r}(x))/3\right)}{4},$ (2)
@@
where $F(u_{r}(x))$ denotes the flatness. In the inertial range, for $\eta\ll r\ll L$ (with $\eta=(\nu^{3}/\varepsilon)^{\frac{1}{4}}$, being the Kolmogorov length scale), the scale dependence is expected to follow the relation kolmogorov1962refinement; castaing1990velocity,
$\Lambda^{2}(r)=\Lambda_{0}^{2}+\mu/9\;\ln(L/r),$} (3)
@@
with $\Lambda_{0}^{2}$ a constant that depends on the properties of the flow (see also figure 1). Note, that the relation (3) is consistent with the power-law scaling behavior of the energy spectral density in the inertial range, as discussed below (see equation (4)).
@@
Within this rationale, the non-Gaussianity of the velocity increment distributions can be quantified by a single scalar parameter. The value of $\mu$ has been commonly accepted to be constant at large values of $Re_{\lambda}$. The reported values range between $0.2$ and $0.4$ for HIT and a value of $0.26$ has been established as representative of turbulent flows arneodo1996structure; vinnes2023characterizing; neunaber2020distinct.
Despite intense research on turbulent flows, no relation between $\mu$ and $C_{\varepsilon}$ has been predicted nor found. Furthermore, the universality of these constants remains an open question, with a growing amount of evidence pointing against this feature. While most studies deal with homogeneous turbulence, in this study we also focus on inhomogeneous or non-ideal turbulent flows. The data selection is explained below. To analyse such non-ideal and/or non-HIT data, we use equation (1) for $C_{\varepsilon}$ and equations 2 and 3 for $\mu$. For $C_{\varepsilon}$, we justify our approach by considering each velocity time series as a 1 D-surrogate. This was done before, for example for fractal-grid turbulence vassilicos2015dissipation and turbulent boundary layer flows nedic2017dissipation. For the estimation of $\mu$, we justify this approach based on the quality of the fits of the probability functions, $p(u_{r}(x))$, as shown in figure 9 in the appendix SM. Our approach aims to extend our understanding of ideal turbulent flow conditions. In particular, we aim to quantify the intermittency of such flows in a general way, independently of the turbulence model used to interpret the results. Therefore, a good fit of $p(u_{r}(x))$ is sufficient for our objectives. It is important to mention that the refined estimation of $\mu$ used in this work does not require an assumption regarding the behavior of the structure functions.
Hence, we present a detailed experimental work on turbulent wakes supported by previously published datasets on grid-generated turbulence and a turbulent axisymmetric jet.
@@
Figure 1: Schematic representation of the spatial expansion of the turbulence cascade: the blue curve shows the energy spectrum, with energy transported to smaller scales, while the red curve illustrates the evolution of the shape factor $\Lambda^{2}(r)$ for normalized velocity increments. Note that in this sketch, $\Lambda_{0}^{2}$ is approximately zero.
I Experimental setup
@@
We conducted wind tunnel experiments in LEGI, Grenoble by generating a turbulent wake behind porous and impermeable objects under different inflow conditions. The facility is a closed-circuit system with test section dimensions of 0.75 $\times$ 0.75 $\times$ 44\, m3 (see the appendix SM). The objects were cylinders and disks while both the diameter and solidity vary (see also table I in the appendix SM). To reduce the blockage, the cylinders were directly screwed to the walls of the tunnel while the disks were hold at the tunnel’s centreline with four thin piano wires which went through taped holes in the tunnel walls. For varying the inflow, both the entrance of the test section and the tunnel freestream velocity were modified. We used laminar inflow, and different turbulent background flows generated by a static regular grid with cylindrical bars or an active grid driven in a triple random mode (information about the turbulent flow produced by grids is reported in a previous work ferran2023experimental). The wake of the objects was scanned by a 11\,D-hot-wire with a temporal resolution of 5050\,kHz (with an anti-aliasing filter set at 3030\,kHz) which was traversed in streamwise and spanwise directions. The conversion from the temporal to the spatial domain was made using Taylor’s hypothesis. The nearest streamwise position to the wake generator was 5d5\,d and the farthest was 200d200\,d, with $d$ being the diameter of the object. For every position the data was acquired for 120120\,s. In total, $15$ configurations were carried out and $498$ points were taken. For simplicity, points in the neighborhood of the centreline are considered, including regions with stronger shear (pre-analysis shows that our results also hold for non-centreline data).
Additionally, the background flows without the object were also measured (see table I in the appendix SM for further information.)
@@
As discussed above, we complement our study with $110$ velocity time series from three previously published experimental datasets covering different canonical flows ferran2023characterising; mora2019energy; renner2001experimental (for further details see the appendix SM). Within those measurement configurations, five involve decaying grid-generated turbulence and one captures a free axisymmetric jet. All cases use air as the fluid and the velocity time signals were collected by means of a 11\,D-hot-wire.
II Results
@@
For all those points from all mentioned datasets, a rigorous analysis of common 11\,D-free-shear turbulence constants is derived and checked individually. Only time series that have the following characteristics were kept for further analysis:
@@
a) the turbulence intensity $TI=u^{\prime}/\overline{u}$ with $\overline{u}$ being the actual mean velocity, decreases in streamwise direction, b) the large scale increments have an almost Gaussian distribution ($\Lambda_{0}^{2}<0.025$) c) the slope of the energy spectrum within the inertial range is smaller than 1.25-1.25, d) the length of the inertial range for the energy spectrum is larger than $0.5$ decades, e) the $R^{2}$-value of the linear regression of the fit of $\Lambda^{2}$ is larger than $0.99$, f) the mean squared deviation from the calculation of $\Lambda^{2}$ is in average smaller than $0.11$, g) the flatness of the velocity time signal is around 33 ($2.67<{\color[rgb]{0,0,0}F_{u}}<3.1$) and h) the skewness of the time signal is below $0.1$. Details and an example of the analysis is shown and discussed in the appendix SM. After this conditioning, 312 velocity time series remained for the analysis. Each of them represents at least $2\times 10^{4}$ integral time scales.
@@
The integral length scale $L$ is calculated by integrating the autocorrelation function until it crosses $1/\mathrm{e}$ ($\mathrm{e}$ being the Eulerian number). Other integration limits were tested, and the trends presented in this Letter remained unchanged. $\varepsilon$ is derived – the common definition for HIT conditions is used – by integrating the dissipation spectrum $\varepsilon=\int_{0}^{\infty}15\nu\>k^{2}E(k)\>\mathrm{d}k$ (with $k$ the wavenumber in the Fourier space) and modeling the high-frequency noise with an extrapolated fit (a test with different fitting limits and functions gave an estimated error in the estimation of $\varepsilon$ in the order of $5\,\%$). We extracted, after equation (3), $\mu$ by fitting the evolution of the shape factor $\Lambda^{2}$ within the inertial range castaing1990velocity.
@@
First, we investigate if the parameters $\mu$ and $C_{\varepsilon}$ present a constant value in our dataset, including all the flows studied. In figure 2a) it is clear that $\mu$ presents a large spreading and it does not approach towards an asymptotic limit at large values of $Re_{\lambda}$. (For reference, the dependency of $\mu$ with the streamwise distance to the turbulence generator is shown in the appendix SM). Figure 2b) shows that $C_{\varepsilon}$ also behaves similarly, i.e. $C_{\varepsilon}$ does not have a constant value.
@@
Having a closer look on the data we see that for each set of boundary conditions, a linear relation between $C_{\varepsilon}$ and $\sqrt{Re_{G}}/Re_{\lambda}$ is consistent with the dissipation scalings from non-equilibrium turbulence vassilicos2015dissipation. Remarkably, when such conditions are changed, the linear relation still applies but with a different slope, showing the emergence of new length scales that affect the dissipation scalings.
@@
Overall, our dataset does not show a universal or predictable behaviour in terms of neither $C_{\varepsilon}$ nor $\mu$. Next, we want to see if there is a relation between $\mu$ and $C_{\varepsilon}$. Figure 3a) shows that $\mu\propto 1/C_{\varepsilon}$, or, respectively, that the product of those two quantities is constant. In order to quantify this, initially, a simple fit is presented through the data points. Note that the commonly accepted value of $\mu$ of around 0.26 is not the asymptotic limit of our measurements. Moreover, the product of $\mu$ and $C_{\varepsilon}$, here defined as $\alpha$, is shown as a function of $Re_{\lambda}$ in figure 3b). Remarkably, $\alpha$ not only remains approximately constant for all cases, it also does not show any trends with $Re_{\lambda}$. Moreover, the well-centered PDF of $\alpha$ values indicates a Gaussian distribution. In the appendix we provide further figures and a table showing that the value of the standard deviation of $\alpha$ remains lower than the individual corresponding values of the standard deviation of $C_{\varepsilon}$ and $\mu$ SM. To the authors best knowledge, this is the first time a clear relation between these constants is found.
(a)                             (b)
Figure 2: a) $\mu$ as a function of $Re_{\lambda}$ for the highly restricted data. The red line corresponds to a commonly accepted value for $\mu$ for homogeneous isotropic turbulence arneodo1996structure. b) $C_{\varepsilon}$ versus $\sqrt{Re_{G}}/Re_{\lambda}$ for the highly restricted data. In general, the symbols in the legend are identical for all figures throughout this manuscript and correspond to the configurations in table I in the appendix SM. For laminar inflow, squared markers are used. For the regular grid and the active grid, markers are shaped as circles and triangles, respectively. For cylinders as generators, the markers contains a “//” (case names start with “C”) while for disks a $\backslash$” (case names start with “D”) is used. Hollow circular markers means no object and is equivalent to grid turbulence (case names start with “G”). The “x” marker indicates a free jet (case names start with “J”). $N_{v}$ indicates the number of velocity time series shown in this plot.
(a)                             (b)
@@
Figure 3: a) $\mu$ as a function of $C_{\varepsilon}$ for the highly restricted data. The red line indicates a commonly accepted value for $\mu$ for homogeneous isotropic turbulence arneodo1996structure, and the black line corresponds to a least-square fit ($R^{2}=0.86$ with the fitted values of $0.106\pm 0.002$ and $0.022\pm 0.009$). b) $\alpha$ versus $Re_{\lambda}$ for the highly restricted data with the PDF of $\alpha$ values $p(\alpha)$. The red line indicates the result for $\alpha$ from the fit from a) while the black solid line represents the actual mean value of the ensemble of $\alpha$ values. Additionally, two black dashed lines indicate the corresponding standard deviations $\sigma$ from the mean. For both figures, the symbols and corresponding configurations are shown and explained in figure 2 and table I in the appendix SM. $N_{v}$ indicates the number of velocity time series shown in this plot.
@@
We continue by investigating how two other relevant parameters behave, namely the Kolmogorov constant $C_{k}$ and the power law governing the energy spectral density within the inertial range. Following the notation of mydlarski1996onset,
$\color[rgb]{0,0,0}{E(k)=C_{k}\>\>\varepsilon^{2/3}\>k^{-5/3}\>(k\>\eta/2\>\pi)^{-\gamma+5/3},}$} (4)
the K41 prediction for the spectra is complemented with the correction $(k\>\eta/2\>\pi)^{-\gamma+5/3}$. K41 is therefore recovered when $\gamma=5/3$ and, for large values of $Re_{\lambda}$, $C_{k}$ is a constant (commonly accepted to be around 0.5 for HIT sreenivasan1995universality). The value of 5/3 emerges through a dimensional argument which followed the first two similarity hypotheses of Kolmogorov. Over the last years, several works reported values of $\gamma$ that deviate from the expected 5/3 value steiros2022balanced; rodriguez2023not; neunaber2020distinct, even at large values of $Re_{\lambda}$, most likely due to intermittency corrections. On the other hand, less evidence of deviations from $C_{k}\,\approx\,0.5$ has been reported in the literature meyers2008functional).
@@
In figure 4a) & b), the dependency of $\gamma$ on $C_{\varepsilon}$ as well as $C_{k}$ on $\gamma$ are presented, respectively ($C_{k}$ as a function of $C_{\varepsilon}$ is shown in the appendix SM). Interestingly, the relation between $\mu$ and $C_{\varepsilon}$ is transferred to the other quantities. Moreover, all quantities present large variations with respect to the expected values for HIT within the K41 phenomenology. We remark that our results are still consistent with those reported in other experiments in inhomogeneous flows, such as other turbulent wakes neunaber2020distinct or atmospheric turbulence rodriguez2023not.
(a)                             (b)
@@
Figure 4: a) $\gamma$ as a function of $C_{\varepsilon}$ and b) $C_{k}$ versus $\gamma$ for the highly restricted data. The red lines indicate both the commonly accepted value for $C_{k}$ and $\gamma$ for HIT sreenivasan1995universality, while the black lines correspond to a least-square fit for a) and b) with $R^{2}$ being $0.8$ and $0.99$, respectively. Once more, the symbols and corresponding configurations are shown and explained in figure 2 and table I in the appendix SM. $N_{v}$ indicates the number of velocity time series shown in this plot.
III Discussion